# Sven Matthiesen2,   Rainer Stiefelhagen1

### v1

@@ Abstract
As Vision-Language Models (VLMs) advance, Assistive Technologies (ATs) for helping People with Visual Impairments (PVIs) are evolving into generalists, capable of performing multiple tasks simultaneously. However, benchmarking VLMs for ATs remains under-explored in the literature. To bridge this gap, we first create a novel AT benchmark (@Bench). Guided by a pre-design user study with PVIs, our benchmark includes the five most crucial vision-language tasks: Panoptic Segmentation, Depth Estimation, Optical Character Recognition (OCR), Image Captioning, and Visual Question Answering (VQA). Additionally, we propose a novel AT model (@Model) that addresses all tasks simultaneously and can be expanded to more assistive functions for helping PVIs. Our framework exhibits outstanding performance across tasks by integrating multi-modal information, and it offers PVIs a more comprehensive scene understanding. Extensive experiments prove the effectiveness and generalizability of our framework.
@@ 2.1 Assistive Technologies for the Blind
A common goal in ATs is to develop artificial intelligent systems via vision-language algorithms to help PVIs. VizWiz, introduced by Bigham et al. in 2010 [3], presents the first multi-task datasets and artificial intelligence challenges originating from PVIs. These include over 10 tasks, such as VQA [10], image captioning [11], object detection [33] and object classification [2], etc. It covers various scenes in the daily life of PVIs and provides valuable data for the research of ATs. Other task-specific datasets [37, 47] focus on the recognition of obstacles and tactile paving, further contributing to this field. Based on the PVIs-oriented and general datasets, most work used visual model to address the daily challenges encountered by PVIs. For example, previous works [1, 39, 46, 9] employ visual tasks such as detection, segmentation and depth estimation to accomplish avoidance, navigation, privacy protection, etc. Some other works have focused on language-modal or cross-modal tasks, such as OCR [30, 25], image captioning [11], and VQA [10]. Compared to these existing methods, we conduct a human-in-the-loop study to design a unified multi-modal benchmark for evaluating VLMs for ATs.
@@ 2.2 Generalist Vision-Language Models
Generalist VLMs have witnessed remarkable advancements, driven by breakthroughs in deep learning techniques [31, 42, 51]. Developing a generalist model for multiple tasks poses unique challenges due to the heterogeneous inputs and outputs, including RGB images, depth maps, binary masks, bounding boxes, language, etc. Previous methods MetaLM [12] and PaLI [5] use language models as general-purpose interfaces to various foundation models. GLIPv2 [45] unifies both localization and VL understanding tasks as grounded vision-language tasks. OFA [42] and Unified-IO [24] introduce a Seq2Seq framework for the unification of I/O, architectures, tasks, and modalities. X-Decoder [51] can predict pixel-level segmentation and language tokens through a generalized decoding model. Uni-Perceiver v2 [22] formulates different tasks as a unified maximum likelihood estimation problem without any task-specific fine-tuning. However, these methods are either unable to perform multi-task training, or they focus too much on language-modal tasks or cross-modal tasks, while ignoring vision-modal tasks that are important for PVIs, such as segmentation and depth estimation. Therefore, we propose a new method that can comprehensively consider and balance multiple assistance-related uni-modal and cross-modal tasks, and can use one suit of weights for all tasks.
@@ 5.5 Ablation Study
Tokenizer Vocabulary IC13 IC15 SVT IIIT5K SVTP CUTE avg
@@ 6 Discussion
Additionally, we analyze the comments given by the blind user and list a few. About OCR, “Usability by completely blind people would be very important to me. Existing systems of this type have difficulty selecting the right target. Text recognition on a document in front of me or on packaging in my hand is also very possible with a smartphone. However, if the function is able to read door signs, hanging posters or street signs that are not within direct reach, that would be an extremely useful function.”. According to the scores in the questionnaire and comments, PVIs attach great importance to text recognition, especially in some special scenarios. Therefore, unlike other generalist models that ignore OCR task, @Bench contains OCR task and introduce multiple OCR datasets, covering various text scenarios. About object recognition, “This is a function I would definitely expect!”, “It would be important, on the one hand, to have a high level of reliability of recognition and, on the other hand, to be able to determine, even for completely blind people, that the correct object is being recognized.”. By introducing multi-category ADE20K and combining panoptic segmentation, @Model can recognize a large number of stuffs and things, compared with detection, semantic segmentation. At the same time, the more categories of objects are recognized, the accuracy of recognition will increase and the recognition will be more reliable. In sum, these positive comments and great suggestions inspire our work and provide guidance for future work of exploring VLMs for assistive technology.
@@ 7 Conclusion
Acknowledgement. This work was supported in part by the Ministry of Science, Research and the Arts of Baden-Wurttemberg (MWK) through the Cooperative Graduate School Accessibility through AI-based Assistive Technology (KATE) under Grant BW6-03, in part by the Federal Ministry of Education and Research (BMBF) through a fellowship within the IFI program of the German Academic Exchange Service (DAAD), and in part by Future Mobility Grants from InnovationCampus Future Mobility (ICM). We thank HoreKA@KIT, HAICORE@KIT, and bwHPC supercomputer partitions.
@@ B.1 Pixel-level Output Loss
Segmentation Loss. There are two losses on the segmentation corresponding to two tasks. For mask classification, we use text encoder $\mathbf{Enc_{T}}$ to encode all $N$ class names including “background” into $N$ text embeddings $\mathbf{E}_{cls}{\in}\mathcal{R}^{N\times C}$ and take it to represent the concept. Afterward, we take the first ($m-1$) latent queries and compute the dot-product between these outputs and concept embeddings to obtain an affinity matrix $\mathbf{S}_{cls}{\in}\mathcal{R}^{(m-1){\times}N}$ and compute $\mathcal{L}_{cls}{=}\mathbf{CE}(\mathbf{S}_{cls},\mathbf{y}_{cls})$, with the ground-truth class $\mathbf{y}_{cls}$. For mask prediction, we use Hungarian matching [4, 6] to find the matched entries of first ($m-1$) outputs to ground-truth annotations. Afterward, we to use binary cross-entropy loss $\mathcal{L}_{bce}$ and dice loss $\mathcal{L}_{dice}$ to compute the loss. Thus, the overall training loss function of panoptic segmentation is:
@@ D.1 Comments on Generalist Assistance Systems
By conducting the questionnaire survey, we communicated with visually impaired individuals to comprehend the functionalities they expect a generalist assistive system should possess. We got some thoughts like: “It should find the door, look for stairs in an open area, read the house/room number, read signs/plates, describe the environment, warn me of obstacles, and can navigate the corridors with a floor plan.”. Some participants also described specific usage scenarios, “I would use navigation and obstacle detection systems outside. It should warn me of obstacles or describe something I’m about to encounter. For example, if I’m navigating outside and there’s a road ahead, then it should say if it has a roundabout or an intersection. Or, if there is a railroad crossing, announce something similar. It would be cool if there was an all around view. The system says, the front of you is street and the back is a building, left is bike racks, etc. If there is a name of the store, read it out. The most important thing is to have a general navigation ability based on this all around view. If I then say navigate to the store (name of the store) recognized by this system, then it should navigate me there.”. Based on these thoughts and comments, essential functions identified by People with Visual Impairments (PVIs) that a generalist assistive system should include are:
@@
    Comprehensive Scene Interpretation. PVIs expressed a desire for a system that provides a holistic view of their surroundings. This “all-around vision” function should describe streets, buildings, and other elements in the vicinity.
@@ D.2 More Comments
Navigation. The majority of participants (5 P: 5 participants) prioritize outdoor navigation, noting its greater complexity and risk. They highlight that outdoor environments pose larger obstacles, longer and more complex routes, and a higher likelihood of getting lost compared to indoor scenarios. One participant emphasized, “Outdoor navigation is much more important. Indoors, the reach of a cane is much more likely to adequately capture the surroundings. The distances are shorter and the density of people is higher.”. Another added, “Definitely outdoors. If I have to go into a building I don’t know, it will probably only be for once. It’s not worth learning a way to do that.”. The unpredictability of outdoor spaces, such as traffic, was also mentioned as a significant factor. Conversely, a minority (2 P) believes that indoor navigation is more important. They mention the challenges of navigating within large unfamiliar buildings, locating specific rooms, stairs, elevators, or exits, walking across a large open-area, and walking in rooms with highly differentiated structures, such as restrooms. Importantly, they spend most of their time indoors.
@@
Text Recognition. Today, PVIs mainly use screen readers to recognize digital texts and usually use smartphones, or smartphones Apps to read non-digital texts. However, they find non-digital text reading is difficult and cumbersome, like “Everything I receive on paper in the post annoys me. I use apps like Seeing AI and Be My Eyes or the iPhone’s magnifying glass to read non-digital texts, but using a smart glass to read these text directly would be better.”. They also pointed out that it is also important for them to read signs to find the right floor or hallway and read door numbers to enter the right room.
@@
Other Functions. About depth estimation, “This function helps one develop a mental map of an environment. You get the proportions well.”. About object location, “In my personal environment I am always very sure where all the things I am looking for are. However, locating a true one in a larger shelf section of 3-4 meters would be very useful. A function that detects objects that don’t belong in that space would also be very helpful to check a room for overlooked clutter. The glasses could use a reference photo of the tidy room and then report any anomalies, such as dirty dishes on the table or socks on the floor.” and “If I only need it if I can’t find something in my apartment, it could make the search easier, but I would need it pretty rarely.”. About surroundings understanding, “It would be important to me that the description be highly efficient. The short form is always first” and “Most of the time we are not interested in the scene because it is too much information for us. But descriptions of photos, environments, etc. are very exciting. ChatGPT is really great.”. About scene recognition, “Perhaps interesting for recognize different scenarios, but a correspondingly efficient description of the image would serve the same purpose. I can’t imagine a situation where I would need room detection. I usually know which room I’m going to or being led into.”. About visual Q&A, “This function would make it possible to expand a short initial description of an image dynamically and according to your own needs. That would improve the overall function enormously.”.
@@
Table 7: Comparison of results of mixed training for different tasks. Note:“Other”, “Unanswerable”, “Yes/No”, “Number” are 4 different answer types for VQA. (PS = panoptic segmentation, IC = image captioning).
@@ E.1 Complementariness in Multitasking
As shown in the experiments section, our @Model exhibits a strong performance in captioning and VQA under multi-task training. Here, we further study the role of segmentation objectives in vision-language (VL) understanding, as well as the role of different vision-language understanding tasks on each other. To investigate, we mix different tasks for training. In Table 7, for captioning, when jointly trained with VQA or PS, or all tasks, CIDEr improved by $1.2$, $1.7$ and $7.4$ respectively. For VQA, we report 5 numbers for better analysis, namely the accuracy for 4 Q&A types: other, unanswerable, yes/no, number, and the overall accuracy. From the comparison of these numbers, when training VQA alone, the model tends to predict “unanswerable” to improve the accuracy. Because in the dataset, the unanswerable type of Q&A is the most common. For other types of Q&A, the accuracy is relatively lower because a deeper or more granular understanding of the semantic information of image is required to predict the correct answers. After joint training with captioning, the accuracy of unanswerable type Q&A decreased, and the accuracy of other types increased. The model does not just return “unanswerable” blindly but understands more semantic information of the image and then make predictions. When all tasks are trained together, the accuracy of other type Q&A is greatly improved ($+8.9\%$). We analyze that it is because the question of this type of Q&A is usually “what is this?”, and the segmentation task naturally has a very good assisting effect in answering this question. Segmentation data can help models to learn more fine-grained visual understanding and consequently benefit vision-language tasks. We also give some examples to show these improvements in Fig. 7. Along with our findings in the main paper, we conclude that segmentation has clear benefits to VL learning and different VL tasks are complementary to each other.

### v2

@@ Abstract
As Vision-Language Models (VLMs) advance, human-centered Assistive Technologies (ATs) for helping People with Visual Impairments (PVIs) are evolving into generalists, capable of performing multiple tasks simultaneously. However, benchmarking VLMs for ATs remains under-explored. To bridge this gap, we first create a novel AT benchmark (@Bench). Guided by a pre-design user study with PVIs, our benchmark includes the five most crucial vision-language tasks: Panoptic Segmentation, Depth Estimation, Optical Character Recognition (OCR), Image Captioning, and Visual Question Answering (VQA). Besides, we propose a novel AT model (@Model) that addresses all tasks simultaneously and can be expanded to more assistive functions for helping PVIs. Our framework exhibits outstanding performance across tasks by integrating multi-modal information, and it offers PVIs a more comprehensive assistance. Extensive experiments prove the effectiveness and generalizability of our framework.
@@ 2.1 Assistive Technologies for the Blind
A common goal in ATs is to develop artificial intelligent systems via vision-language algorithms to help PVIs. VizWiz, introduced by Bigham et al. in 2010 [3], presents the first multi-task datasets and artificial intelligence challenges originating from PVIs. These include over 10 tasks, such as VQA [10], image captioning [11], object detection [35] and object classification [2], etc. It covers various scenes in the daily life of PVIs and provides valuable data for the research of ATs. Other task-specific datasets [39, 49] focus on the recognition of obstacles and tactile paving, further contributing to this field. Based on the PVIs-oriented and general datasets, most work used visual model to address the daily challenges encountered by PVIs. For example, previous works [1, 41, 48, 9, 51, 24] employ visual tasks such as detection, segmentation and depth estimation to accomplish avoidance, navigation, privacy protection, etc. Some other works have focused on language-modal or cross-modal tasks, such as OCR [32, 27], image captioning [11], and VQA [10]. Compared to these existing methods, we conduct a human-in-the-loop study to design a unified multi-modal benchmark for evaluating VLMs for ATs.
@@ 2.2 Generalist Vision-Language Models
Generalist VLMs have witnessed remarkable advancements, driven by breakthroughs in deep learning techniques [33, 44, 55]. Developing a generalist model for multiple tasks poses unique challenges due to the heterogeneous inputs and outputs, including RGB images, depth maps, binary masks, bounding boxes, language, etc. Previous methods MetaLM [12] and PaLI [5] use language models as general-purpose interfaces to various foundation models. GLIPv2 [47] unifies both localization and VL understanding tasks as grounded vision-language tasks. OFA [44] and Unified-IO [26] introduce a Seq2Seq framework for the unification of I/O, architectures, tasks, and modalities. X-Decoder [55] can predict pixel-level segmentation and language tokens through a generalized decoding model. Uni-Perceiver v2 [22] formulates different tasks as a unified maximum likelihood estimation problem without any task-specific fine-tuning. However, these methods are either unable to perform multi-task training, or they focus too much on language-modal tasks or cross-modal tasks, while ignoring vision-modal tasks that are important for PVIs, such as segmentation [50] and depth estimation [23]. Therefore, we propose a new method that can comprehensively consider and balance multiple assistance-related uni-modal and cross-modal tasks, and can use one suit of weights for all tasks.
@@ 5.5 Ablation Study
Tokenizer Vocabulary IC13 IC15 SVT IIIT5K SVTP CUTE Avg.
@@ 6 Discussion
Additionally, we analyze the comments given by the blind user and list a few. About OCR, “Usability by completely blind people would be very important to me. Existing systems of this type have difficulty selecting the right target. Text recognition on a document in front of me or on packaging in my hand is also very possible with a smartphone. However, if the function is able to read door signs, hanging posters or street signs that are not within direct reach, that would be an extremely useful function.". According to the scores in the questionnaire and comments, PVIs attach great importance to text recognition, especially in some special scenarios. Therefore, unlike other generalist models that ignore OCR task, @Bench contains OCR task and introduce multiple OCR datasets, covering various text scenarios. About object recognition, “This is a function I would definitely expect!", “It would be important, on the one hand, to have a high level of reliability of recognition and, on the other hand, to be able to determine, even for completely blind people, that the correct object is being recognized.". By introducing multi-category ADE20K and combining panoptic segmentation, @Model can recognize a large number of stuffs and things, compared with detection, semantic segmentation. At the same time, the more categories of objects are recognized, the accuracy of recognition will increase and the recognition will be more reliable. In sum, these positive comments and great suggestions inspire our work and provide guidance for future work of exploring VLMs for assistive technology.
@@ 7 Conclusion
Acknowledgment. This work was supported in part by the Ministry of Science, Research and the Arts of Baden-Wurttemberg (MWK) through the Cooperative Graduate School Accessibility through AI-based Assistive Technology (KATE) under Grant BW6-03, in part by the Federal Ministry of Education and Research (BMBF) through a fellowship within the IFI program of the German Academic Exchange Service (DAAD), and in part by Future Mobility Grants from InnovationCampus Future Mobility (ICM). We thank HoreKA@KIT, HAICORE@KIT, and bwHPC supercomputer partitions.
@@ B.1 Pixel-level Output Loss
Segmentation Loss. There are two losses on the segmentation corresponding to two tasks. For mask classification, we use text encoder $\mathbf{Enc_{T}}$ to encode all $N$ class names including “background" into $N$ text embeddings $\mathbf{E}_{cls}{\in}\mathcal{R}^{N\times C}$ and take it to represent the concept. Afterward, we take the first ($m-1$) latent queries and compute the dot-product between these outputs and concept embeddings to obtain an affinity matrix $\mathbf{S}_{cls}{\in}\mathcal{R}^{(m-1){\times}N}$ and compute $\mathcal{L}_{cls}{=}\mathbf{CE}(\mathbf{S}_{cls},\mathbf{y}_{cls})$, with the ground-truth class $\mathbf{y}_{cls}$. For mask prediction, we use Hungarian matching [4, 6] to find the matched entries of first ($m-1$) outputs to ground-truth annotations. Afterward, we to use binary cross-entropy loss $\mathcal{L}_{bce}$ and dice loss $\mathcal{L}_{dice}$ to compute the loss. Thus, the overall training loss function of panoptic segmentation is:
@@ D.1 Comments on Generalist Assistance Systems
By conducting the questionnaire survey, we communicated with visually impaired individuals to comprehend the functionalities they expect a generalist assistive system should possess. We got some thoughts like: “It should find the door, look for stairs in an open area, read the house/room number, read signs/plates, describe the environment, warn me of obstacles, and can navigate the corridors with a floor plan.". Some participants also described specific usage scenarios, “I would use navigation and obstacle detection systems outside. It should warn me of obstacles or describe something I’m about to encounter. For example, if I’m navigating outside and there’s a road ahead, then it should say if it has a roundabout or an intersection. Or, if there is a railroad crossing, announce something similar. It would be cool if there was an all around view. The system says, the front of you is street and the back is a building, left is bike racks, etc. If there is a name of the store, read it out. The most important thing is to have a general navigation ability based on this all around view. If I then say navigate to the store (name of the store) recognized by this system, then it should navigate me there.". Based on these thoughts and comments, essential functions identified by People with Visual Impairments (PVIs) that a generalist assistive system should include are:
@@
    Comprehensive Scene Interpretation. PVIs expressed a desire for a system that provides a holistic view of their surroundings. This “all-around vision" function should describe streets, buildings, and other elements in the vicinity.
@@ D.2 More Comments
Navigation. The majority of participants (5 P: 5 participants) prioritize outdoor navigation, noting its greater complexity and risk. They highlight that outdoor environments pose larger obstacles, longer and more complex routes, and a higher likelihood of getting lost compared to indoor scenarios. One participant emphasized, “Outdoor navigation is much more important. Indoors, the reach of a cane is much more likely to adequately capture the surroundings. The distances are shorter and the density of people is higher.". Another added, “Definitely outdoors. If I have to go into a building I don’t know, it will probably only be for once. It’s not worth learning a way to do that.". The unpredictability of outdoor spaces, such as traffic, was also mentioned as a significant factor. Conversely, a minority (2 P) believes that indoor navigation is more important. They mention the challenges of navigating within large unfamiliar buildings, locating specific rooms, stairs, elevators, or exits, walking across a large open-area, and walking in rooms with highly differentiated structures, such as restrooms. Importantly, they spend most of their time indoors.
@@
Text Recognition. Today, PVIs mainly use screen readers to recognize digital texts and usually use smartphones, or smartphones Apps to read non-digital texts. However, they find non-digital text reading is difficult and cumbersome, like “Everything I receive on paper in the post annoys me. I use apps like Seeing AI and Be My Eyes or the iPhone’s magnifying glass to read non-digital texts, but using a smart glass to read these text directly would be better.". They also pointed out that it is also important for them to read signs to find the right floor or hallway and read door numbers to enter the right room.
@@
Other Functions. About depth estimation, “This function helps one develop a mental map of an environment. You get the proportions well.". About object location, “In my personal environment I am always very sure where all the things I am looking for are. However, locating a true one in a larger shelf section of 3-4 meters would be very useful. A function that detects objects that don’t belong in that space would also be very helpful to check a room for overlooked clutter. The glasses could use a reference photo of the tidy room and then report any anomalies, such as dirty dishes on the table or socks on the floor." and “If I only need it if I can’t find something in my apartment, it could make the search easier, but I would need it pretty rarely.". About surroundings understanding, “It would be important to me that the description be highly efficient. The short form is always first" and “Most of the time we are not interested in the scene because it is too much information for us. But descriptions of photos, environments, etc. are very exciting. ChatGPT is really great.". About scene recognition, “Perhaps interesting for recognize different scenarios, but a correspondingly efficient description of the image would serve the same purpose. I can’t imagine a situation where I would need room detection. I usually know which room I’m going to or being led into.". About visual Q&A, “This function would make it possible to expand a short initial description of an image dynamically and according to your own needs. That would improve the overall function enormously.".
@@
Table 7: Comparison of results of mixed training for different tasks. Note:“Other", “Unanswerable", “Yes/No", “Number" are 4 different answer types for VQA. (PS = panoptic segmentation, IC = image captioning).
@@ E.1 Complementariness in Multitasking
As shown in the experiments section, our @Model exhibits a strong performance in captioning and VQA under multi-task training. Here, we further study the role of segmentation objectives in vision-language (VL) understanding, as well as the role of different vision-language understanding tasks on each other. To investigate, we mix different tasks for training. In Table 7, for captioning, when jointly trained with VQA or PS, or all tasks, CIDEr improved by $1.2$, $1.7$ and $7.4$ respectively. For VQA, we report 5 numbers for better analysis, namely the accuracy for 4 Q&A types: other, unanswerable, yes/no, number, and the overall accuracy. From the comparison of these numbers, when training VQA alone, the model tends to predict “unanswerable" to improve the accuracy. Because in the dataset, the unanswerable type of Q&A is the most common. For other types of Q&A, the accuracy is relatively lower because a deeper or more granular understanding of the semantic information of image is required to predict the correct answers. After joint training with captioning, the accuracy of unanswerable type Q&A decreased, and the accuracy of other types increased. The model does not just return “unanswerable" blindly but understands more semantic information of the image and then make predictions. When all tasks are trained together, the accuracy of other type Q&A is greatly improved ($+8.9\%$). We analyze that it is because the question of this type of Q&A is usually “what is this?”, and the segmentation task naturally has a very good assisting effect in answering this question. Segmentation data can help models to learn more fine-grained visual understanding and consequently benefit vision-language tasks. We also give some examples to show these improvements in Fig. 7. Along with our findings in the main paper, we conclude that segmentation has clear benefits to VL learning and different VL tasks are complementary to each other.